OpenStudy (anonymous):

Will fan solve 5.4+0.2x=-1.4x+8.6 a.2 b.-2 c..2 d.20

OpenStudy (anonymous):

You are cool

OpenStudy (anonymous):

@BleuSoul57 thanks

OpenStudy (anonymous):

mop

OpenStudy (camerondoherty):

\[5.4+0.2x=-1.4x+8.6\]Thats ur equation, right?

OpenStudy (anonymous):

@camerondoherty yes

OpenStudy (camerondoherty):

Ok, so first you add 1.4x to both sides Can you do that?

OpenStudy (anonymous):

5.4+1.6+8.6?

OpenStudy (camerondoherty):

You mean? \[5.4+1.6x=8.6 \]

OpenStudy (camerondoherty):

@BleuSoul57

OpenStudy (anonymous):

yes

OpenStudy (camerondoherty):

Ok, so now, you subtract 5.4 from both sides c:

OpenStudy (anonymous):

1.6x=3.2

OpenStudy (anonymous):

now you need to get x by itself

OpenStudy (anonymous):

Divide 1.6 by 1.6 and do the same to the opposite side \[ \frac{1.6x }{ 1.6 }=\frac{ 3.2 }{ 1.6 } \] That leaves x alone on the left side and the number 2 one the right side \[x=2 \] To check redo the whole equation with your new information (x=2) Check: 5.4+1.6(2)=8.6